[Brahms-l] Shift report 20050217 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Feb 17 2005 - 00:00:02 EST
Shifter leader was hy
Shifters were hhd


1. Access until 9:00pm, followed by machine setup. Got beam for physics at around 11:15pm. But they kept it for only half an hour.
2. Pedestal runs: 13207, 13208;
   Setting Fs4A3450 Mrs40A2255
   RunNo   Trig2    Trig3
   13212   94k       46k 
(they dumped the beam for unexpected beam abort)
3. Hiro and Bob installed Left SMX; They are working well, see e-log for  
    info in detail


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 17 00:00:38 2005

This archive was generated by hypermail 2.1.8 : Thu Feb 17 2005 - 00:00:51 EST